### v1

@@ Abstract.
The complexity of the Generally Accepted Accounting Principles (GAAP) and the hierarchical structure of eXtensible Business Reporting Language (XBRL) filings make financial auditing increasingly difficult to automate and verify. While large language models (LLMs) have demonstrated strong capabilities in unstructured text understanding, their ability to reason over structured, interdependent, and taxonomy-driven financial documents remains largely unexplored. To fill this gap, we introduce 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first taxonomy-aligned, structure-aware, multi-document benchmark for evaluating LLMs on financial auditing tasks. Built from real US-GAAP-compliant XBRL filings, 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: defines three complementary subtasks, FinSM for semantic consistency, FinRE for relational consistency, and FinMR for numerical consistency, each targeting a distinct aspect of structured auditing reasoning. We further propose a unified evaluation framework integrating retrieval, classification, and reasoning metrics across these subtasks. Extensive zero-shot experiments on 13 state-of-the-art LLMs reveal that current models perform inconsistently across semantic, relational, and mathematical dimensions, with accuracy drops of up to 60–90% when reasoning over hierarchical multi-document structures. Our findings expose the systematic limitations of modern LLMs in taxonomy-grounded financial reasoning and establish 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: as a foundation for developing trustworthy, structure-aware, and regulation-aligned financial intelligence systems. The benchmark dataset is available at Hugging Face [^1] FinAuditing.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Applied computing Extensible Markup Language (XML)$c$s: Information systems Test collectionsccs: Information systems Information extractionccs: Information systems Question answering
Figure 1. Trends in the proportion of financial restatements among US companies from 2014 to 2024, categorized by reissuance and revision restatements. Data adapted from the Financial Times report, ”Accounting errors force US companies to pull statements in record numbers” (Dec 9, 2024).
@@ 1. Introduction
The Generally Accepted Accounting Principles (GAAP) provide the foundation for corporate financial reporting in the United States, but their growing complexity and continuous updates have made compliance verification increasingly difficult (Debreceny et al., 2010; Hoitash and Hoitash, 2018). Financial restatements caused by misapplied GAAP rules have become a persistent source of reputational and regulatory risk, with recent reports indicating that severe reissuance restatements now account for the majority of filing corrections as shown in Figure. 1. While the introduction of the eXtensible Business Reporting Language (XBRL) has standardized digital disclosures, its hierarchical and interdependent structure introduces new challenges for automated auditing and anomaly detection.
With the rapid advancement of large language models (LLMs), researchers have increasingly explored their applicability across a wide range of tasks in the financial domain, including information extraction (IE) (Loukas et al., 2022; Sharma et al., 2023; Wang et al., 2025), text generation (TG) (Mukherjee et al., 2022; Xie et al., 2023), text analysis (TA) (Sinha and Khandait, 2021), question answering (QA) (Chen et al., 2021, 2022; Zhu et al., 2021; Zhao et al., 2022, 2023), risk management (RM) (Feng et al., 2023), and forecasting (FO) (Soun et al., 2022), showing potential to improve GAAP interpretation and understanding in XBRL filings. However, as summarized in Table 1, existing benchmarks mainly rely on unstructured or semi-structured data drawn from a single document or limited contexts, such as financial reports and news, while overlooking inter-dependencies, compositional hierarchies, and relational constraints that characterize real-world XBRL documents. Moreover, most financial benchmarks assess LLMs on isolated understanding and reasoning abilities, focusing on surface-level semantic understanding within plain texts or single documents. They rarely evaluate deeper capabilities such as structured semantic retrieval, hierarchical relation understanding, and multi-step reasoning, which are fundamental for capturing the structural and logical complexity inherent in financial data. Furthermore, while benchmarks such as FinTagging (Wang et al., 2025) initially utilize financial taxonomies, most do not fully leverage their hierarchical and relational structures, leaving taxonomy-driven reasoning largely unexplored. This omission limits the ability of current evaluations to measure how well LLMs can reason over structured, taxonomy-driven, and cross-document dependencies, which are crucial for a comprehensive understanding of finance.
@@
Table 1. Comparison of existing financial NLP benchmarks and 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. The table shows whether each benchmark supports structured retrieval (Struct.Retrieval), hierarchical information extraction (Hiera.IE), multi-step numerical reasoning (MultiStep.Reasoning), taxonomy-driven supervision (Need.Taxonomy), and at the cross-document level (At.Cross-Doc).
Benchmark Source Data Format Task Struct.Retrieval? Hiera.IE? MultiStep.Reasoning? Need.Taxonomy? At.Cross-Doc?
FiNER (Loukas et al., 2022) Financial reports Text IE
FNXL (Sharma et al., 2023) Financial reports Text IE
Headlines (Sinha and Khandait, 2021) Financial news Text TA
ECTSum (Mukherjee et al., 2022) Earnings call transcripts Text TG
CCF (Feng et al., 2023) Credit card transactions Text RM
BIGDATA22 (Soun et al., 2022) Tweets, historical prices Text FO
FinTagging (Wang et al., 2025) Financial reports Text & Table IE, IR
FinQA (Chen et al., 2021) Financial reports Text & Table QA
ConvFinQA (Chen et al., 2022) Financial reports Text & Table QA
TAT-QA (Zhu et al., 2021) Financial reports Text & Table QA
MultiHiertt (Zhao et al., 2022) Financial reports Semi-structured and hierarchical Text&Table QA
DOCMATH-EVAL (Zhao et al., 2023) Financial reports Semi-structured document segmentation QA
FinAuditing (ours) XBRL filings Structured and hierarchical multi-document segmentation IE, IR, QA
To address these gaps, we propose 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first benchmark that evaluates LLMs on structured semantic retrieval, hierarchical relation understanding, and multi-step reasoning over taxonomy-driven and interconnected XBRL filings, focusing on detecting inconsistencies and errors across structured financial documents. To this end, 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: encompasses three independent tasks: the financial semantic matching (FinSM) task, which examines whether models can detect semantic inconsistencies between financial disclosures and standardized taxonomy concepts; the financial relationship extraction (FinRE) task, which evaluates the ability to identify structural and hierarchical errors among related financial elements; and the financial mathematical reasoning (FinMR) task, which measures whether models can reason over numerical relationships to verify the logical correctness of reported values. To comprehensively assess these tasks, we design task-specific evaluation metrics that collectively capture retrieval accuracy, structural understanding, and numerical reasoning performance, using measures such as Hit Rate, Macro-F1, and error-rate indicators for a multi-perspective evaluation of LLM capabilities. As a result, our work introduces a more comprehensive and systematic evaluation framework that measures LLMs’ capabilities on structured, hierarchical, and multiple financial documents.
We conduct an extensive evaluation on 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: under the zero-shot setting, covering 13 state-of-the-art (SOTA) proprietary and open-source LLMs across three structured auditing tasks. Our results show that while leading models such as DeepSeek-V3 in FinSM, GPT-4o in FinRE, and Fin-o1-14B in FinMR achieve the highest performance within their respective tasks, all models exhibit substantial limitations. DeepSeek-V3 demonstrates better retrieval accuracy yet struggles with coverage and semantic alignment across interconnected structured documents. GPT-4o shows strong hierarchical relation understanding across all three relation types, whereas most open-source models have difficulty with the “CombinationErr” type. Fin-o1-14B, despite its domain specialization, remains weak in schema interpretation and structural consistency during numerical reasoning. Overall, these findings highlight that even frontier models fall short of reliable structured auditing, revealing persistent challenges in semantic retrieval, relational reasoning, and multi-step computation over hierarchical XBRL data.
We conclude our main contributions as follows:
@@
[…]
@@ 2. Related work
Existing financial NLP research has made progress in information extraction (Loukas et al., 2022; Sharma et al., 2023; Wang et al., 2025), semantic understanding (Mukherjee et al., 2022; Xie et al., 2023; Sinha and Khandait, 2021), and numerical reasoning (Chen et al., 2021, 2022; Zhu et al., 2021; Zhao et al., 2022, 2023), yet most benchmarks focus on unstructured or semi-structured data from single documents, such as financial reports or news articles, overlooking the structured and interconnected nature of XBRL filings.
Early work on XBRL mainly targeted localized tasks. Loukas et al. (Loukas et al., 2022) developed an XBRL tagging benchmark for numerical entity extraction, and Sharma et al. (Sharma et al., 2023) proposed numeral labeling within XBRL sentences. Recently, Wang et al. (Wang et al., 2025) explored structured tag normalization via LLMs based on the US-GAAP taxonomy. Steven et al. (Katz et al., 2024) employed ChatGPT for tag retrieval, while Han et al. (Han et al., 2024) introduced XBRL-Agent to assess LLMs’ analytical and numerical abilities on XBRL data. However, these studies evaluate narrow subtasks and fail to capture the hierarchical and relational dependencies that define real-world XBRL structures.
Beyond XBRL, benchmarks such as FinQA (Chen et al., 2021), ConvFinQA (Chen et al., 2022), TAT-QA (Zhu et al., 2021), MultiHiertt (Zhao et al., 2022), and DocMath-Eval (Zhao et al., 2023) assess reasoning over textual and tabular data, yet they remain text-centric and lack taxonomy-driven supervision. To fill this gap, our proposed 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides the first benchmark to evaluate LLMs on structured, hierarchical, and cross-document reasoning grounded in real XBRL filings.
@@
Detecting errors across the hierarchical and structured financial documents requires understanding semantics, hierarchical relations, and numerical logic. As shown in Figure 2, to evaluate these capabilities, we propose 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a benchmark that measures how well LLMs identify semantic, structural, and numerical inconsistencies across multi-document XBRL filings.
Figure 2. Overview of the 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark framework for evaluating error detection on XBRL filings across three tasks.
3.1. Preliminary knowledge
The US-GAAP Taxonomy serves as a standardized vocabulary and hierarchical framework that defines both the financial concepts (e.g., us-gaap:CashAndDueFromBanks) and the relationships among them. It provides a consistent semantic structure for financial reporting, ensuring comparability and enabling automated analysis.
An XBRL filing is a structured digital financial report formatted in XML. It consists of multiple interconnected documents, including an instance document, schema document, and several linkbases, which together define and link the representations and relationships of financial elements (see Table 2 for details).
Additionally, Fact corresponds to the reported values associated with US-GAAP tags in the instance documents. Context specifies details about a reported fact (e.g. reporting period). Axis represents a categorical dimension that allows financial value to be classified from different perspectives. For instance, ”RevenueByGeographyAxis” organizes revenue based on geographic regions. Member is a specific value within an Axis, representing a particular category. For example, within ”RevenueByGeographyAxis”, ”USMember” denotes revenue generated in the United States, while ”EuropeMember” refers to European revenue.
Table 2. Main Components of XBRL Filings
Component Description

### v2

@@ Abstract.
Going beyond simple text processing, financial auditing requires detecting semantic, structural, and numerical inconsistencies across large-scale disclosures. As financial reports are filed in XBRL, a structured XML format governed by accounting standards, auditing becomes a structured information extraction and reasoning problem involving concept alignment, taxonomy-defined relations, and cross-document consistency. Although large language models (LLMs) show promise on isolated financial tasks, their capability in professional-grade auditing remains unclear. We introduce 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a taxonomy-aligned, structure-aware benchmark built from real XBRL filings. It contains 1,102 annotated instances averaging over 33k tokens and defines three tasks: Financial Semantic Matching (FinSM), Financial Relationship Extraction (FinRE), and Financial Mathematical Reasoning (FinMR). Evaluations of 13 state-of-the-art LLMs reveal substantial gaps in concept retrieval, taxonomy-aware relation modeling, and consistent cross-document reasoning. These findings highlight the need for realistic, structure-aware benchmarks. We release the evaluation code [^1] https://github.com/The-FinAI/FinAuditing and dataset [^2] https://huggingface.co/collections/TheFinAI/finauditing publicly, and the task currently serves as the official benchmark of an ongoing public evaluation contest [^3] https://open-finance-lab.github.io/SecureFinAI_Contest_2026/.
@@
$c$s: Applied computing Extensible Markup Language (XML)$c$s: Information systems Test collectionsccs: Information systems Information extractionccs: Information systems Question answering
Figure 1. An illustrative example of cross-document financial inconsistency.
@@ 1. Introduction
Consider a scenario where a global corporation reports “Total Cash” as \$100M in its high-level summary, while detailed supporting disclosures across sub-documents sum to only \$95M. Detecting such inconsistencies is a non-trivial information verification problem. In modern financial reporting, this information spans thousands of pages of structured filings, making manual verification nearly impossible. While ensuring data accuracy is vital for market transparency, identifying these “hidden” inconsistencies remains challenging even for advanced AI systems, due to the massive scale and structural complexity involved.
Modern financial disclosures are governed by Generally Accepted Accounting Principles (GAAP) and operationalized via the eXtensible Business Reporting Language (XBRL) (Debreceny et al., 2010). Unlike plain text, these filings are heterogeneous, hierarchical, and highly interdependent (Debreceny et al., 2010; Hoitash and Hoitash, 2018). Auditing these documents requires more than keyword search, as it demands sophisticated structured information retrieval (SIR) and reasoning (Li and Zhai, 2015; da Silva, 2018). Models must semantically align concepts across disparate tables, interpret nested hierarchical metadata, and maintain logical consistency across multiple interdependent documents. When these capabilities fall short, the consequences are tangible: firms issue restatements, i.e., costly corrections of previous reports, which have risen persistently as disclosure complexity outpaces current analytical tools [^4] https://www.ft.com/content/716c4ad5-e8fa-4a34-afba-9fb2d1db019d.
@@
Table 1. Comparison of existing financial NLP benchmarks and 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. We compare support for structured retrieval (Struct.Retrieval), hierarchical information extraction (Hiera.IE), multi-step numerical reasoning (MultiStep.Reasoning), taxonomy-driven supervision (Need.Taxonomy), and cross-document reasoning (At.Cross-Doc).
Benchmark Struct.Retr. Hiera.IE MultiStep.Reas. Need.Taxon. At.Cross-Doc
FiNER (Loukas et al., 2022)
FNXL (Sharma et al., 2023)
FinTagging (Wang et al., 2025)
FinQA (Chen et al., 2021)
ConvFinQA (Chen et al., 2022)
TAT-QA (Zhu et al., 2021)
MultiHiertt (Zhao et al., 2022)
DOCMATH-EVAL (Zhao et al., 2023)
@@
Recent advances in large language models (LLMs) have sparked interest in applying them to financial tasks, such as information extraction (Loukas et al., 2022; Sharma et al., 2023; Wang et al., 2025), question answering (Chen et al., 2021, 2022; Zhu et al., 2021; Zhao et al., 2022, 2023), and numerical reasoning (Soun et al., 2022) and trading (Qian et al., 2025a; Li et al., 2025). However, it remains unclear whether these models can support professional-grade auditing, which requires consistent reasoning over hierarchical structures, taxonomy-defined relations, and cross-document dependencies. This raises a critical question: how should we evaluate models on structured, regulation-driven financial information tasks? As summarized in Table 1, existing financial NLP benchmarks exhibit three fundamental limitations: (1) Contextual Isolation: Most benchmarks operate on isolated text snippets or tables, failing to capture the cross-document dependencies that are central to real financial filings (Zhao et al., 2022). (2) Surface-Level Evaluation: Prior work predominantly evaluates local semantic or numerical correctness, without assessing multi-step structural reasoning over reporting hierarchies (Zhao et al., 2023). (3) Taxonomy Underutilization: Although official GAAP taxonomies define valid financial concepts and relations, few works evaluate whether models can retrieve and reason over these constraints (Wang et al., 2025).
To address these gaps, we introduce 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a taxonomy-aligned, structure-aware, multi-document benchmark derived from real-world US-GAAP-compliant XBRL filings. 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: reframes financial auditing as a structured information reasoning problem, and defines three complementary subtasks: Financial Semantic Matching (FinSM) for concept-level semantic alignment, Financial Relationship Extraction (FinRE) for structural and relational consistency, and Financial Mathematical Reasoning (FinMR) for taxonomy-grounded numerical verification. Together, these tasks shift evaluation from isolated text understanding to holistic reasoning over structured financial information, providing a high-fidelity testbed for financial capabilities under formal constraints. Our main contributions are threefold:
  • A Taxonomy-Grounded Financial Resource: We release the first benchmark built from real-world XBRL filings that explicitly integrates multi-document structures and official GAAP taxonomies for evaluation.
  • Structure-Aware Task Design: We propose three complementary subtasks (FinSM, FinRE, FinMR) that capture semantic, relational, and numerical reasoning challenges inherent to structured financial information.
  • Systematic Evaluation of Intelligent Systems: We provide an initial evaluation of 13 SOTA LLMs, while emphasizing that 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is model-agnostic and designed to support the evaluation of complex architectures, including retrieval-augmented generation (RAG) pipelines and multi-agent auditing systems, for structured financial information retrieval and reasoning.
@@ 2. Related work
Prior financial NLP research has advanced information extraction (Loukas et al., 2022; Sharma et al., 2023; Wang et al., 2025), semantic understanding (Mukherjee et al., 2022; Xie et al., 2023; Sinha and Khandait, 2021), and numerical reasoning (Chen et al., 2021, 2022; Zhu et al., 2021; Zhao et al., 2022, 2023). However, most benchmarks focus on unstructured or semi-structured data from individual documents, such as financial reports or news articles, and overlook the structured, hierarchical, and interconnected nature of XBRL filings. Early XBRL-focused studies primarily address localized subtasks, including numerical entity extraction (Loukas et al., 2022), sentence-level numeral labeling (Sharma et al., 2023), tag normalization (Wang et al., 2025), and retrieval- or agent-based evaluation of LLMs on XBRL data (Katz et al., 2024; Han et al., 2024). While informative, these efforts evaluate narrow components in isolation and do not capture the taxonomy-level dependencies that characterize real-world XBRL structures. Related benchmarks on tabular and textual reasoning, such as FinQA (Chen et al., 2021), ConvFinQA (Chen et al., 2022), TAT-QA (Zhu et al., 2021), MultiHiertt (Zhao et al., 2022), and DocMath-Eval (Zhao et al., 2023), similarly lack taxonomy-driven supervision. In contrast, 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is designed to evaluate LLMs on structured, hierarchical, and cross-document reasoning grounded in realistic XBRLs.
Figure 2. Overview of the 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: construction pipeline.
@@
Detecting errors in structured financial documents requires semantic, structural, and numerical reasoning. As shown in Figure 2, we introduce 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a benchmark for evaluating LLMs in identifying inconsistencies across multi-document XBRL filings.